40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jun 3 16:46:31 2019, gautam, Update, IOO, IMC not locking PCdrive_RMS.png
    Reply  Tue Jun 4 10:56:31 2019, gautam, Update, IOO, IMC diagnostics IMCdiag.pdf
       Reply  Thu Jun 6 22:11:53 2019, Koji, Update, IOO, IMC diagnostics 
          Reply  Thu Jun 13 18:01:18 2019, aaron, Update, IOO, IMC diagnostics 
             Reply  Fri Jun 14 13:10:00 2019, aaron, Update, IOO, IMC diagnostics 
                Reply  Mon Jun 17 12:37:16 2019, aaron, Update, IOO, IMC diagnostics 
                   Reply  Tue Jun 18 20:35:07 2019, aaron, Update, IOO, IMC diagnostics 
                      Reply  Wed Jun 19 19:12:51 2019, aaron, Update, IOO, IMC diagnostics PMC_Error_Spectrum.pdfPDH_signal.pdfPDH_signal_full.pdf
                         Reply  Mon Jun 24 15:49:05 2019, aaron, Update, IOO, IMC diagnostics sweep_config_updated.png
                            Reply  Tue Jul 2 19:36:18 2019, aaron, Update, IOO, IMC diagnostics PDH_error.pdfPMC_Error_Spectrum.pdf
          Reply  Fri Jun 21 19:36:26 2019, Koji, Update, IOO, IMC diagnostics IMC_error.pdf
             Reply  Sun Jun 23 08:09:53 2019, gautam, Update, IOO, NPRO diagnostics consolidatedOscopeScreenCaps.pdfRINcomp.pdfPLL_OLTF.pdfPLLnoise.pdf
                Reply  Sun Jun 23 09:36:32 2019, gautam, Update, IOO, IMC is locking normally again PMCdiag.pdf
                   Reply  Sun Jun 23 14:43:14 2019, Koji, Update, IOO, IMC is locking normally again 
                   Reply  Mon Jun 24 08:12:10 2019, gautam, Update, IOO, IMC is locking normally again IMCdutycycle.png
                      Reply  Mon Jun 24 11:48:35 2019, gautam, Update, IOO, IMC in-loop error spectra and OLTF IMCfreqNoise.pdfIMC_OLTF.pdfIMC_CMboard.jpg
Message ID: 14691     Entry time: Mon Jun 24 11:48:35 2019     In reply to: 14690
Author: gautam 
Type: Update 
Category: IOO 
Subject: IMC in-loop error spectra and OLTF 

Attachment #1 - In loop error spectra, measured as Koji posted end of last week.

  • Main difference is that the line noise seems much lower.
  • For the "dark" measurement, I set the IN1 gain of the servo board to the value of +4 dB, which is what it is in lock.
  • As Koji mentioned, this isn't an apple-to-apple comparison as the IMC loop will squish the plotted orange trace.
  • Nevertheless, the fact that the blue trace is above orange everywhere gives confidence that we are in fact measuring frequency noise.
  • For the higher frequency measurement, I used the AG4395 analyzer, which has 50 ohm input impedance. So to get the measurements with the SR785 to line up, I multiplied these by x2.
  • For the frequency axis calibration, I used the value of 13 kHz/V for the PDH discriminant, which was what I measured it to be last year (but I didn't check again today).
  • Note that the IMC locking loop OLTF has not been undone, so this isn't the actual laser frequency noise on the transmitted beam. In order to measure the latter, we'd have to use (for example) an arm cavity as an analyzer.

Attachment #2 - OLTF of the IMC loop.

  • Measurement was made using the IN1/IN2 method, injection was done at the "A EXC" front panel BNC input.
  • For comparison, I've overlaid a measurement from the 2017 IMC loop investigations. Doesn't seem to be significantly different.
  • UGF and phase margin are in the ballpark of what they were reported to be in the past.

Attachment #3 - Photo courtesy Koji showing the bank of BNC connectors used for these measurements.

Clearly, these measurements were taken in a time when the IMC was "well behaved". How to characterize what's happening when this isn't the case?

Attachment 1: IMCfreqNoise.pdf  73 kB  Uploaded Mon Jun 24 12:55:09 2019  | Hide | Hide all
IMCfreqNoise.pdf
Attachment 2: IMC_OLTF.pdf  73 kB  Uploaded Mon Jun 24 12:55:17 2019  | Hide | Hide all
IMC_OLTF.pdf
Attachment 3: IMC_CMboard.jpg  193 kB  Uploaded Mon Jun 24 12:59:24 2019  | Hide | Hide all
IMC_CMboard.jpg
ELOG V3.1.3-